OpenStudy (anonymous):

6 4√4 – 7 3√4 + 2 3√4 – 3 4√4 a. -2√4 b. -2√8 c. 3 4√4- 5 3√4 d. 3 4√4- 5 3√8

OpenStudy (mayankdevnani):

64*2-73*2+23*2-34*2 128-146+46-68=-40

OpenStudy (anonymous):

=6 4√4 – 7 3√4 + 2 3√4 – 3 4√4 =24 √4 - 21 √4 + 6 √4 - 12 √4 = -3 √4 So the answer is c, because of 3 4√4- 5 3√4 = -3 √4 = 12√4- 15√4 = -3 √4